# On the Moduli Space of Discrete Morse Functions

### v1

@@ 2. $\Leftarrow$
Let $\varphi\colon X\rightarrow Y$ be either a simplicial map between regular CW complexes or a non-degenerate map between CW complexes. We define an induced map $\varphi_{*}\colon\mathbb{R}^{X}\rightarrow\mathbb{R}^{Y}$ on the spaces of discrete Morse functions inductively as follows:
@@
    $\varphi(f)_{\tau}\coloneqq f_{\sigma}$
@@
    Start with the minimal dimensional such $\tau$s.
    • If $\operatorname{Face}_{1}(\tau)=\emptyset$ and $\operatorname{Coface}_{1}(\tau)=\emptyset$, then define $\varphi_{*}(f)_{\tau}\coloneqq 0.$
    • If $\operatorname{Face}_{1}(\tau)=\emptyset$ and $\operatorname{Coface}_{1}(\tau)\neq\emptyset$, then define $\varphi_{*}(f)_{\tau}\coloneqq\operatorname{min}(f_{\lvert\operatorname{Coface}_{1}(\tau)})-1$.
    • If $\operatorname{Face}_{1}(\tau)\neq\emptyset$ and $\operatorname{Coface}_{1}(\tau)=\emptyset$, then define $\varphi_{*}(f)_{\tau}\coloneqq\operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+1$.
@@
      $\varphi_{*}(f)_{\tau}\coloneqq\frac{\operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+\operatorname{up}(f_{\lvert\operatorname{Coface}_{1}(\tau)})}{2}$
      for any $f\in\mathbb{R}^{X}$, where
      $\operatorname{up}(f_{\lvert\operatorname{Coface}_{1}(\tau)})\coloneqq\begin{cases}\operatorname{min}(f_{\lvert\operatorname{Coface}_{1}(\tau)})&\text{ if }f\text{ is defined on some cell of }\operatorname{Coface}_{1}(\tau).\\ \operatorname{min}(f_{\lvert\operatorname{Coface}(\tau)})&\text{ if }f\text{ is only defined on cells of }\operatorname{Coface}(\tau)\setminus\operatorname{Coface}_{1}(\tau)\\ \operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+2&\text{ if }f\text{ is not defined on any cell of }\operatorname{Coface}(\tau).\end{cases}$
      Here, we slightly abuse notation because $f_{\lvert\operatorname{Coface}_{1}(\tau)}$ might not be defined for all cells of $\operatorname{Coface}_{1}(\tau)$. If $f_{\lvert\operatorname{Coface}_{1}(\tau)}$ is only defined for some cells of $\operatorname{Coface}_{1}(\tau)$, we take the minimum on those cells.
@@
The set $\operatorname{Face}_{1}(\sigma)$ corresponds to the set of elements $\tau$ of $D(X)$ which are covered by $\sigma$, whereas the cells of $\operatorname{Coface}_{1}(\sigma)$ are exactly the cells $\tau$ of $X$ which cover $\sigma$. If one uses the covering relation instead of the codimension, there is a straightforward generalization to non-regular CW complexes.
@@
The map $\varphi_{*}$ preserves induced matchings if $\varphi$ is injective.
@@
If $\varphi$ in injective, then $X$ can be interpreted as a subcomplex of $Y$. The map $\varphi_{*}(f)$ is then just an extension of $f$ to $Y$ that is critical on $Y\setminus\varphi(X)$. By construction, we have $\varphi_{*}(f)_{\lvert\varphi(X)}=f\circ\varphi^{-1}$ and, therefore, the statement holds. ∎
@@
We consider the following simplicial map between two regular complexes:
@@
Let $\varphi\colon X\rightarrow Y$ be a cellular map of CW complexes. Then there is an induced map $\varphi^{*}\colon\mathbb{R}^{Y}\rightarrow\mathbb{R}^{X}$ given by $\varphi^{*}g\coloneqq g\circ\varphi$, for $g\colon Y\rightarrow\mathbb{R}$.
@@
Let $X\rightarrow Y$ be an injective non-degenerate map of CW complexes. Then $\varphi^{*{-1}}(\mathcal{A}(X))\subset\mathcal{A}(Y)$ is a subarrangement. Moreover, $\varphi^{*}$ restricts to a map $\mathcal{M}(Y)\rightarrow\mathcal{M}(X)$.
@@
The first statement follows directly by the definition of non-degenerate maps because they preserve cover relations. The second statement follows by injectivity. ∎
4. Relationship to smooth Morse theory
@@
The codimension of $a$ is defined as $\operatorname{codim}(a)\coloneqq\lvert\{q\in f^{-1}(a)|p\text{ critical}\}\rvert-1$.
@@
[…]
@@
such that the natural stratification of $\mathcal{F}$ is the preimage of the stratification induced by the hyperplane arrangement $\mathcal{H}_{q}$ (see Example 2.6).
Our plan for this section is to identify the space $\mathbb{R}^{q}$ in Proposition 4.2 with the space of discrete functions on the CW decomposition of $M$ induced by $f$, to investigate how far the neighborhood $\mathcal{V}$ extends, and to generalize $\eta$ in order to define maps which compare the space of smooth Morse functions on $M$ to spaces of discrete Morse functions on cellular decompositions of $M$.
It is well known (see e.g. [BH04, Theorem 6.34]) that Morse functions only induce a cellular decomposition $M_{f}$ of $M$, unique up to cell equivalence, if they fulfill the Morse–Smale property. Alternatively, on can associate gradient-like Morse–Smale vector fields to non-Morse–Smale functions instead (see [Mat02, Chapter 3 and Section 4.2]). Since we want to take the point of view of the moduli space of Morse functions, we have fixed a Riemannian metric and consider the subspace $\mathcal{MS}_{\text{smo}}(M)\subset\mathcal{M}_{\text{smo}}(M)$ of Morse–Smale functions on $M$.
@@
Let $Z$ be either the space of Morse functions $\mathcal{M}_{\text{smo}}(M)\subset\mathcal{F}$ or the space of Morse–Smale functions $\mathcal{MS}_{smo}(M)\subset\mathcal{M}_{\text{smo}}(M)\subset\mathcal{F}$ and let $f\in Z$. We denote by $\mathcal{N}(f)\subset Z$ the connected component of $f$ in $Z$ [^6] It will be clear from the context which of the two possibilities for $Z$ will be used..
@@
It is a classical result that the property of being a Morse function is stable, i.e. for any Morse function $f$, there is an open neighborhood $U(f)\subset C^{\infty}(M)$ such that $U(f)\subset\mathcal{M}_{\text{smo}}(M)$ (see [BH04, Corollary 5.24]). We want to investigate this phenomenon in a slightly different direction: we show that the CW complex induced by a Morse–Smale function is stable on connected components of Morse–Smale functions. As a first step towards that, we have the following theorem:

### v2

@@ 3. The Space of Discrete Morse Functions
Let $\varphi\colon X\rightarrow Y$ be a non-degenerate map between CW complexes. We define an induced map $\varphi_{*}\colon\mathbb{R}^{X}\rightarrow\mathbb{R}^{Y}$ on the spaces of discrete functions. We define $\varphi_{*}(f)$ for any discrete function $f\in\mathbb{R}^{X}$ as follows:
@@
    $\varphi(f)(\tau)\coloneqq f(\sigma)$
@@
    We define
[…]
    Here, we slightly abuse notation because $f_{\lvert\operatorname{Coface}_{1}(\tau)}$ might not be defined for all cells of $\operatorname{Coface}_{1}(\tau)$. If $f_{\lvert\operatorname{Coface}_{1}(\tau)}$ is only defined for some cells of $\operatorname{Coface}_{1}(\tau)$, we take the minimum on those cells. Start with the minimal dimensional $\tau$ such that $\lvert\varphi^{-1}(\tau)\rvert\neq 1$.
    • If $\operatorname{Face}_{1}(\tau)=\emptyset$ and $\operatorname{Coface}_{1}(\tau)=\emptyset$, then define
      $\varphi_{*}(f)_{\tau}\coloneqq 0.$
    • If $\operatorname{Face}_{1}(\tau)=\emptyset$ and $\operatorname{Coface}_{1}(\tau)\neq\emptyset$, then define
[…]
    • If $\operatorname{Face}_{1}(\tau)\neq\emptyset$ and $\operatorname{Coface}_{1}(\tau)=\emptyset$, then define
      $\varphi_{*}(f)_{\tau}\coloneqq\operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+1.$
@@
      $\varphi_{*}(f)_{\tau}\coloneqq\frac{\operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+\operatorname{up}(f_{\lvert\operatorname{Coface}_{1}(\tau)})}{2}.$
@@
We call $\varphi_{*}$ the push forward map induced by $\varphi$.
@@
The set $\operatorname{Face}_{1}(\sigma)$ corresponds to the set of elements $\tau$ of $F(X)$ which are covered by $\sigma$, whereas the cells of $\operatorname{Coface}_{1}(\sigma)$ are exactly the cells $\tau$ of $X$ which cover $\sigma$. If one uses the covering relation instead of the codimension, there is a straightforward generalization to non-regular CW complexes.
@@
The map $\varphi_{*}$ restricts to a map of spaces of discrete Morse functions $\varphi_{*}\colon\mathcal{M(X)}\rightarrow\mathcal{M}(Y)$. Moreover, $\varphi_{*}$ preserves induced matchings if $\varphi$ is injective.
@@
If $\varphi$ in injective, then $X$ can be interpreted as a subcomplex of $Y$. The map $\varphi_{*}(f)$ is then just an extension of $f$ to $Y$ that is critical on $Y\setminus\varphi(X)$. By construction, we have $\varphi_{*}(f)_{\lvert\varphi(X)}=f\circ\varphi^{-1}$ and, therefore, the second statement holds.
For cells $\tau$ such that $\lvert\varphi^{-1}(\tau)\rvert\neq 1$, we observe that according to Definition 3.31 $\varphi_{*}(f)$ is constructed such that every face that is covered by $\tau$ gets a strictly smaller function value than $\tau$ and every coface that covers $\tau$ gets a strictly greater function value than $\tau$ under $\varphi_{*}(f)$. Hence, every $\tau$ such that $\lvert\varphi^{-1}(\tau)\rvert\neq 1$ is by construction a critical cell of $\varphi_{*}(f)$. In particular, $\varphi_{*}(f)$ is a discrete Morse function. ∎
@@
We consider the following non-degenerate map between two regular complexes:
@@
As a particular important class of maps of CW complexes that induce maps on the spaces of discrete Morse functions, we consider internal collapses and extensions. The case of simple extensions is already covered by injective non-degenerate maps because simple extensions induce inclusions. The case of simple collapses of free faces is a bit more difficult because a pair of a free face and its unique coface in general does not induce a canonical simplicial map. In this case, we only have a (up to homotopy) canonical cellular homotopy equivalence. This still allows us to construct restricted versions of the push forward and the pullback map:
@@
Let $\varphi\colon X\rightarrow X\setminus\{\sigma,\tau\}$ be a simple collapse, where $\sigma$ is the free face of $\tau$ that is being collapsed. Then we define the push forward map induced by $\varphi$ by $\varphi_{*}\colon\mathcal{M}(X)_{|x_{\sigma}\geq x_{\tau}}\rightarrow\mathcal{M}(X\setminus\{\sigma,\tau\}),\ \varphi_{*}(f)(\xi)\coloneqq f(\xi)$ for $\xi\in X\setminus\{\sigma,\tau\}$ and where $\mathcal{M}(X)_{|x_{\sigma}\geq x_{\tau}}$ is the component of $\mathcal{M}(X)$ where $\sigma$ and $\tau$ are matched.
We define the pullback map induced by $\varphi$ by $\varphi^{*}\colon\mathcal{M}(X\setminus\{\sigma,\tau\})\rightarrow\mathcal{M}(X)_{|x_{\sigma}\geq x_{\tau}}$, $\varphi^{*}(f)(\xi)\coloneqq f(\xi)$ for $\xi\in X\setminus\{\sigma,\tau\}$, $\varphi^{*}(f)(\sigma)=\varphi^{*}(f)(\tau)\coloneqq\operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+1$.
Example 3.38.
We consider the following example of a simple collapse.
$\sigma$$\tau\rightarrow$$\varphi$
We illustrate the pullback and push forward map in the two separate examples below.
[…]
We extend these constructions to internal collapses and extensions.
Definition 3.39.
Let $\varphi\colon X\rightarrow Y$ be an internal collapse that collapses a pair of cells $(\sigma,\tau)$ and let $\psi\colon Y\rightarrow X$ be the inverse internal extension.
We define the push forward map $\varphi_{*}\colon\mathcal{M}(X)_{|x_{\sigma}\geq x_{\tau}}\rightarrow\mathcal{M}(Y)$ by $\varphi_{*}(f)(\xi)\coloneqq f(\xi)\text{ for }\xi\neq\sigma,\tau$
We define the pullback map as follows:
$\varphi^{*}(f)(\xi)\coloneqq\begin{cases}f(\xi)&\text{for }\xi\neq\sigma,\tau\\ f(\sigma)=f(\tau)\coloneqq\frac{\operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+\operatorname{up}(f_{\lvert\operatorname{Coface}_{1}(\sigma)})}{2}\end{cases},$
where $\operatorname{up}$ is defined as in Definition 3.31 and $\operatorname{Coface}_{1}$ and $\operatorname{Face}_{1}$ are defined as in Definition 3.30.
For the internal extensions, we define the push forward/pullback as the pullback/push forward of the inverse internal collapse.
Example 3.40.
We consider the following example of an internal collapse.
$\sigma$$\tau\rightarrow$$\varphi$
One instance for the push forward map could be:
[…]
And one instance for the pullback map could be:
[…]
@@
The pullback and push forward maps of internal collapses/extensions from Definition 3.39 are well-defined maps of spaces of discrete Morse functions.
Moreover, the pullback map of an internal collapse, respectively the push forward map of an internal expansion, is an embedding of spaces of discrete Morse functions.
@@
For the pullback map induced by an internal collapse, taking the mean value in the construction ensures that the function values are decreasing along the newly created flow line, except for the pair of the internal collapse, on which $\varphi^{*}(f)$ attains the same value. This implies that the pair, which is being collapsed, is matched and $\varphi^{*}(f)$ is strictly monotone along all other (co-)face relations of $\sigma$ and $\tau$. It follows from $f$ being a discrete Morse function by assumption, that the definition of a discrete Morse function is satisfied at all other face relations as well.
For the push forward map it is essential that we restrict to $\mathcal{M}(X)_{|x_{\sigma}\geq x_{\tau}}$, i.e. the regions of $\mathcal{M}(X)$ where $\sigma$ and $\tau$ are matched. Then it follows similarly as in the previous case, that $\varphi_{*}(f)$ fulfills the definition of a discrete Morse function at every face relation due to $f$ fulfilling it.
For the last statement, we observe that pullbacks of internal collapses and push forwards of internal expansions leave values outside the pair $(\sigma,\tau)$ that is being collpased/expanded unchanged. ∎
Remark 3.42.
The push forward and pullback maps induced by elementary and internal collapses/extensions allow us to relate spaces of discrete Morse functions on simple homotopy equivalent CW complexes with one-another. This way, it may make sense to develop the concept of a “space of discrete Morse functions on a simple homotopy type” as a colimit of a diagram of spaces of discrete Morse functions on all representatives of a simple homotopy class with pullbacks and push forwards of all possible elementary and internal collapses and extensions between them. Due to the difficulties of such infinite diagrams and the combinatorial complexity of simple homotopy types, we leave such ideas for future work.
3.1. The path metric on spaces of discrete Morse functions
In the previous part, we did not explicitly mention how we topologize the spaces of discrete Morse functions. Instead, we rather implied that as a subspace of the finite dimensional vector space of discrete functions, we can use any topology, e.g. the one induced by the Euclidean norm.
In Section 4.1 it will turn out that the Euclidean topology is useful for the relation to the smooth case. However, in applications towards topological data analysis it is often useful to be able to use several different specific metrics that induce the same topology. This provides the flexibility to used metrics adjusted to certain algorithmic approaches while other topological guarantees stay true due to preserving the induced topology.
In this subsection, we induce a specific notion of distance between discrete Morse functions on the same given regular CW complex $X$. This specific metric is meant to take the cost of moving through the hyperplanes of the braid arrangement $\mathcal{H}$ and the Morse arrangement $\mathcal{A}$ into account, and will be useful for the stability result Theorem 4.42 for the induced merge tree.
Remark 3.43.
Let $X$ be a regular CW complex and let $\gamma\colon[0,1]\rightarrow\mathcal{M}(X)$ be a path in the space of discrete Morse functions. The path $\gamma$ cannot simultaneously enter an intersection of hyperplanes $\mathfrak{H}\coloneqq\bigcap\limits_{j}H_{\sigma_{j}}^{\tau_{j}},H_{\sigma_{j}}^{\tau_{j}}\in\mathcal{H}$ and leave another one intersection of hyperplanes $\mathfrak{H}\coloneqq\bigcap\limits_{j^{\prime}}H_{\sigma_{{j^{\prime}}}}^{\tau_{{j^{\prime}}}},H_{\sigma_{j^{\prime}}}^{\tau_{j^{\prime}}}\in\mathcal{H}$ due to entering an intersection of hyperplanes is a closed condition and leaving intersections of hyperplanes is an open condition.
Definition 3.44.
Let $X$ be a regular CW complex, let $\mathcal{M}(X)$ be the space of discrete Morse functions on $X$, and let $f,g\in\mathcal{M}(X)$.
Let $\Gamma\coloneqq\operatorname{Map}([0,1]\rightarrow\mathcal{M}(X))$ be the set of paths $\gamma$ from $f$ to $g$ in the space of discrete Morse functions. For any path $\gamma\in\Gamma$, let $\left(\mathfrak{H}_{i}=\bigcap\limits_{j}H_{\sigma_{i_{j}}}^{\tau_{i_{j}}}\right)$ be the sequence of intersections of hyperplanes that $\gamma$ crosses and let $\bar{t}\coloneqq(t_{l})$ be an ordered sequence of times such that for each intersection of hyperplanes $\mathfrak{H}_{i}$ there is a $t_{l}$ such that $\gamma(t_{l})\in\mathfrak{H}_{i}$ and for each face $F_{i}$ between two consecutive intersections of hyperplanes there is a $t_{l}$ such that $\gamma(t_{l})\in F_{i}$. We call any such sequence $(t_{l})$ and admissible $\gamma$ decomposition of $[0,1]$ and denote the collection of all admissible $\gamma$ decompositions by $T(\gamma)$. We define an ordered sequence $\bar{\gamma}\coloneqq(f\eqqcolon h_{0},h_{1},\dots,h_{l},\dots,h_{k+1}\coloneqq g)$ of discrete Morse functions such that each $h_{i}$ for $1\leq i\leq k$ given by: $h_{l}\coloneqq\gamma(t_{l})$ for $l\leq k$.
We define the path distance on $\mathcal{M}(X)$ as follows:
$d(f,g)\coloneqq\inf\limits_{\gamma\in\Gamma}\inf\limits_{\bar{t}\in T(\gamma)}\sum\limits_{l=0}^{k}\lvert\lvert h_{l}-h_{l+1}\rvert\rvert$,
where $\lvert\lvert\cdot\rvert\rvert$ denotes the Euclidean norm.
Proposition 3.45.
The path distance is a metric.
Proof.
Reflexivity, positivity and symmetry follow directly from the definition. The triangle inequality follows by concatenations of paths. ∎
4. Relationship to other concepts in mathematics
4.1. Smooth Morse Theory and Cerf Theory
@@
The codimension of $a$ is defined as $\operatorname{codim}(a)\coloneqq\lvert\{p\in f^{-1}(a)|p\text{ critical}\}\rvert-1$.
@@
We refer to [SHA93, Chapter 1, §3] and [CER70] for background information on the natural stratification of the space of smooth functions $\mathcal{F}$. Before proceeding, we want to recall from [SHA93, Chapter 1, §3] that $\mathcal{F}$ is a smoothly path connected smooth Fréchet manifold. From now on, all paths will be at least $C^{1}$ and path components will refer to $C^{1}$ path components.
@@
Let $f\in\mathcal{M}_{\text{smo}}(M)$ be a smooth Morse function with $q$ critical points. Given a choice of an ordering of the critical points $c_{1},\dots,c_{q}$, there are open neighborhoods $U_{i}$ with $c_{i}\in U_{i}\subset M$ for which there is an open neighborhood $\mathcal{V}$ with $f\in\mathcal{V}\subset\mathcal{F}$ such that all $\tilde{f}\in\mathcal{V}$ are Morse and have exactly one critical point $\tilde{c}_{i}\in U_{i}$ of the same index as $c_{i}$ for all $1\leq i\leq q$. Moreover, the ordering of the critical points of $f$ defines a topological submersion
@@
such that the restriction of the natural stratification of $\mathcal{F}$ to $\mathcal{V}$ is the preimage of the stratification induced by the hyperplane arrangement $\mathcal{H}_{q}$ (see Example 2.9).
Our plan for this section is to identify the space $\mathbb{R}^{q}$ in Proposition 4.2 with the space of discrete functions on the CW decomposition of $M$ induced by $f$, and use the stability of Morse–Smale functions to extend $\eta$ to path components. This allows for comparison maps from path components of the space of smooth Morse functions on $M$ to spaces of discrete Morse functions on certain cellular decompositions of $M$. For that, we recall the notion of Morse–Smale functions.
Definition 4.3 ([BH04, Definition 6.1]).
A Morse function $f\colon M\rightarrow\mathbb{R}$ on a finite dimensional smooth Riemannian manifold $(M,g)$ is said to satisfy the Morse–Smale transversality condition if and only if the stable manifold of $p$ and unstable manifold of $q$ with respect to $f$ intersect transversally [^13] Recall that two submanifolds $N_{1},N_{2}\subset M$ are said to intersect transversally if at each point of the intersection $p\in N_{1}\cap N_{2}$ the tangent spaces of the submanifolds span the tangent space of the ambient manifold: $T_{p}N_{1}+T_{p}N_{2}=T_{p}M$. for all pairs of critical points $p,q$ of $f$. A Morse function that satisfies the Morse–Smale transversality condition is called a Morse–Smale function.
It is well known (see e.g. [BH04, Theorem 6.34]) that Morse functions only induce an explicit cellular decomposition $M_{f}$ of $M$, unique up to cell equivalence [^14] Recall the Definition 3.21 of cell equivalences., if they fulfill the Morse–Smale property and are compatible with the given Riemannian metric. However, it is proved in [FRA79, Proposition 1.6] that the gradient of any Morse–Smale function can be perturbed to be compatible with the Riemannian metric without leaving its path component in the space of Morse–Smale vector fields, in particular while preserving topological conjugacy. Thus, we have a well-defined CW decomposition of $M$ induced by any Morse–Smale function.
Alternatively, one can associate gradient-like Morse–Smale vector fields to non-Morse–Smale functions instead (see [MAT02, Chapter 3 and Section 4.2]). Since we want to take the point of view of the space of Morse functions, we have fixed a Riemannian metric and consider the subspace $\mathcal{MS}(M)\subset\mathcal{M}_{\text{smo}}(M)$ of Morse–Smale functions on $M$.
@@
Let $Z$ be either the space of Morse functions $\mathcal{M}_{\text{smo}}(M)\subset\mathcal{F}$ or the space of Morse–Smale functions $\mathcal{MS}(M)\subset\mathcal{M}_{\text{smo}}(M)\subset\mathcal{F}$ and let $f\in Z$. We denote by $\mathcal{N}(f)\subset Z$ the path component of $f$ in ZZ [^15] It will be clear from the context which of the two possibilities for $Z$ will be used..
@@
It is a classical result that the property of being a Morse function is stable, i.e. for any Morse function $f$, there is an open neighborhood $U(f)\subset C^{\infty}(M)$ such that $U(f)\subset\mathcal{M}_{\text{smo}}(M)$ and for every path $f_{\_}\colon[0,1]\rightarrow U(f)$ there is an $\epsilon>0$ such that $f_{t}$ is Morse for every $t<\epsilon$ (see [BH04, Corollary 5.24]). This classical result is also used in the proof of Proposition 4.2. Moreover, in [FRA79, §1] the stability of Morse–Smale functions is stated in an even stronger way, which we use for the following theorem:
Definition 4.5.
Let $\phi,\phi^{\prime}$ be flows on a closed orientable Riemannian manifold $M$. $\phi$ and $\phi^{\prime}$ are called topologically conjugate if there exists a homeomorphism $M\rightarrow M$ that carries orbits of $\phi$ to orbits of $\phi^{\prime}$ and preserves their orientation.
Theorem 4.6 ([PAL69, Theorem 1.12], reformulation from [FRA79, §1]).
Let $\phi$ be a Morse–Smale flow on a closed orientable Riemannian manifold $M$ and let $\phi^{\prime}$ be a sufficiently small $C^{1}$ perturbation of $\phi$, then $\phi$ and $\phi^{\prime}$ are topologically conjugate.
For the proof of this theorem, we refer to [PAL69, Theorem 1.12].
Theorem 4.7.
Let $f,g$ be Morse–Smale functions on $M$ which are contained in the same path component, i.e. $g\in\mathcal{N}(f)\subset\mathcal{MS}(M)$. Then there is a bijection $\psi\colon\operatorname{Cr}(f)\xrightarrow[]{\cong}\operatorname{Cr}(g)$ such that there is an orbit of $\phi_{f}$ connecting two critical points $c_{1},c_{2}\in\operatorname{Cr}(f)$ if and only if there is an orbit of $\phi_{g}$ connecting $\psi(c_{1}),\psi(c_{2})\in\operatorname{Cr}(g)$.
Proof.
This result is a consequence of the structural stability of Morse–Smale functions, respectively Morse–Smale vector fields, respectively Morse–Smale flows. Let $f_{t}$ be a path from $f$ to $g$ in $\mathcal{MS}(M)$. We choose an open cover on $I$ that consists of such open neighborhoods $U(t_{i})$ of time steps $t_{i}$ that the restriction of $f_{t}$ to $U(t_{i})$ is a small enough $C^{1}$ perturbation so that stability in the sense of [FRA79] holds. Due to $I$ being compact, there is a finite subcover of neighborhoods where the restriction of $f_{t}$ induces topologically conjugate flows. Thus, $f_{t}$ induces topologically conjugate flows for all $t$. Hence, the statement follows. ∎
Corollary 4.8.
For any compact finite-dimensional Riemannian manifold $M$, all Morse–Smale functions which are in the same path component of $\mathcal{MS}(M)$ induce cell-equivalent CW decompositions $M_{f}$ of $M$. That is, there is an up to cell equivalence well-defined CW decomposition $M_{\mathcal{N}}$ associated to any path component $\mathcal{N}$ of the space of Morse–Smale functions on $M$.
Corollary 4.9.
It follows from Theorem 4.7 that Cerf’s map from Proposition 4.2 extends to a map $\eta\colon\mathcal{N}\rightarrow\mathcal{M}(M_{\mathcal{N}})$, where $\mathcal{M}(M_{\mathcal{N}})$ denotes the CW decomposition of $M$ induced by any Morse function in the path component $\mathcal{N}$ of $\mathcal{M}(M)$. Moreover, Cerf’s proof of Proposition 4.2 shows that this instance of $\eta$ is a topological submersion and compatible with the respective stratifications, too.
Building on Corollary 4.8, we can extend the map $\eta$ from Proposition 4.2 to the path component $\mathcal{N}(f)$ of a given Morse–Smale function $f$. By Proposition 4.2, it still follows that $\eta\colon\mathcal{N}(f)\rightarrow\mathcal{M}(M_{\mathcal{N}(f)})\coloneqq\mathcal{M}(M_{f})$ is a topological submersion which is compatible with the respective stratifications.
Using Cerf’s stratification of the space of Morse functions, especially the connection between different path components given by cancellations and creations of pairs of critical points, we want to further extend $\eta$ to certain systems of path components:
Definition 4.10.
Consider the set $\pi_{0}(\mathcal{MS}(M))$ of path components of the space of Morse–Smale functions on $M$. Let $\mathcal{N}_{1},\mathcal{N}_{2}\in\pi_{0}(\mathcal{MS}(M))$ be path components. Then we define a partial order $\leq$ on $\pi_{0}(\mathcal{MS}(M))$ generated by $\mathcal{N}_{1}<\mathcal{N}_{2}$ if for any $f\in\mathcal{N}_{2}$ there is a $g\in\mathcal{N}_{1}$ such that $g$ arises from $f$ by a cancellation of a pair of critical points. We call $\leq$ the cancellation order on $\pi_{0}(\mathcal{MS}(M))$.
Remark 4.11.
It is shown in [CER70, Section 3] that if two generic Morse functions, i.e. elements of $\mathcal{F}^{0}$, are related to each other by a cancellation of one pair of critical points, then all paths in $\mathcal{F}$ between them are homotopic in $\mathcal{F}$ to one which traverses the first stratum $\mathcal{F}^{1}$ in one isolated point and is otherwise entirely contained in $\mathcal{F}^{0}$. Hence, we can think about two path components $\mathcal{N}_{1},\mathcal{N}_{2}$ such that $\mathcal{N}_{1}<\mathcal{N}_{2}$ as being neighbors, only separated by a codimension one stratum which corresponds to a hyperplane in $\mathcal{M}(\mathcal{N}_{2})$ via $\eta$.